OpenStudy (anonymous):

Find the length of the base of a square pyramid if the volume is 48 cubic inches and has a height of 9 inches.

OpenStudy (mathstudent55):

Do you know the formula for the volume of a pyramid?

OpenStudy (anonymous):

no i forgot

OpenStudy (mathstudent55):

Here it is: \(A_{pyramid} = \dfrac{1}{3}Bh\) where B = area of the base

OpenStudy (anonymous):

these are the answers i can choose from: 4 inches 8 inches 16 inches 24 inches

OpenStudy (mathstudent55):

In this case, the base is a square. If the side of the square has length s, the area of the base is \(s^2\). Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@LiveLaughDie

OpenStudy (mathstudent55):

Now enter all the info we have in the formula. The side is unknow, s, and the area of the base is \(s^2\). The volume is 48 in^3. The height is 9 in.

OpenStudy (anonymous):

im horrible at this

OpenStudy (anonymous):

@tkhunny

OpenStudy (mathstudent55):

\(A_{pyramid} = \dfrac{1}{3} Bh\) \(48.~in^3 = \dfrac{1}{3} \times s^2 \times 9 ~in.\) Since all units are inches and cubic inches, I'll leave out the units to simplify the problem. Our answer will be in inches. \(48 = \dfrac{1}{3} \times s^2 \times 9\) On the right side, 9/3 = 3, so we have: \(48 = 3s^2\) Divide both sides by 3: \(16 = s^2\) \(s^2 = 16\) \(s = 4\) The length of the side is 4 in.

OpenStudy (anonymous):

thanks u rock :D :D :D